Find three consecutive integers whose sum is 123.
If the first number is x, what expression would you use to represent the LARGEST number in this equation? (use lower case letters and no spaces).

Answer by richard1234(7193)

If x is the smallest number, then x+1 is the middle, and x+2 is the largest integer. These three add up to 123, so

x + (x+1) + (x+2) = 123

3x + 3 = 123

x = 40

Therefore the three integers are 40, 41, 42 in which 40 + 41 + 42 = 123.